# Vid2Robot: End-to-end Video-conditioned Policy Learning with Cross-Attention Transformers

### v1

@@ III-A2 How well do video-conditioned policies perform when they are shown a task in an unseen video?
III-B Cross-object motion transfer
TABLE II: Cross-object motion transfer success.
@@ III-B Cross-object motion transfer
Figure 7: Qualitative results for cross-object motion transfer. Given a prompt video of placing coke can upright, we rollout the policy with a green can, chips bag, stapler and a soft toy in front of the robot. We observe that our model can infer the motion of place upright in the prompt video and apply it on other objects. There is an implicit notion of pragmatics in the policy as shown by the selection of green can over other objects.
Our policy and baseline were trained with paired videos as discussed in §II-B. This implies that the training data included only those scenarios where the interaction object shown in prompt is present in the current robot observations. But what if we provided a prompt video of one object and tested on other objects. Does it do the same motion as shown in the prompt video? Interestingly, we found our model to perform learned manipulation actions on objects that it has not seen in train set. We call this emergent behavior as cross-object motion transfer.
We compare Vid2Robot with BC-Z for cross object motion transfer ability with five prompt videos, namely, ‘knock water bottle over’, ‘pick green rice chip bag’, ‘place coke can upright’, ‘pick coke can from bottom drawer and place on counter’, and ‘place apple into top drawer’. Each prompt video is evaluated with unrelated objects in robot’s initial observation. The objects used for evaluation are ‘orange’, ‘green can’, ‘chips bag’, ‘banana’, ‘pink piggy soft toy’, ‘wrist watch’. We selected objects to have diverse shape, size, and deformability to evaluate situations that require different grasps for success.
The evaluation setup is similar to §III-A. Here the evaluator sets up one of the object for a task and records rollouts for each model. We compare 2 models on 5 tasks with 6 objects, so every evaluator runs $2\!\times\!5\!\times\!6\!=\!60$ rollouts. We repeat the evaluation with four raters, thus reporting results in Table II on a total of $4\!\times\!60=240$ rollouts.
III-B1 Can we provide a prompt video of one object and test it on other objects? Does the policy do the same motion as shown in the prompt video?
In Fig 7, we show the above experimental setup qualitatively. We use a prompt video to ‘place coke can upright’. We observe that the policy is able transfer the action of ‘placing upright’ to several objects, like a green can, a chips bag, a stapler, and a soft toy. Note that the policy adheres to the prompt video and chooses green can over chips bag or banana for placing upright.
Quantitatively, we observe that BC-Z is often unable to successfully complete the tasks when testing cross=object motion transfer, as shown in each task in Table II. In contrast, our model (34%) performs better than BC-Z (17%) in this setting and performs the motion indicated in the prompt video. Our model is comparable to BC-Z with 45% success rate on picking out-of-distribution objects. More importantly, tasks involving placing into drawers demonstrates significant improvement ($29\%\rightarrow 54\%$). For certain tasks like picking from drawers and placing on counters and knocking over, BC-Z is unable to perform at all whereas Vid2Robot is able to complete the task $25\%-29\%$ of the time.
III-C Ablations
In §II-E, we presented action prediction loss and three auxiliary losses. Here we analyze the role of these additional loss functions to the overall success rate. We investigate the impact of (1) not using any auxiliary loss, and (2) adding auxiliary language loss.
We consider the tasks similar to that described in §III-A, that is, 9 tasks for evaluating each policy.
@@ III-C Ablations
We have 3 model variants, namely, the original Vid2Robot, the one without video-text contrastive loss (CL) and the one with only action prediction loss. We ask 3 human evaluators to run the each model variant with 2 rollouts each. In total, we report results with $3\!\times\!3\!\times\!9\!\times\!2\!=\!162$ rollouts in Fig 8. The error bars indicate the standard deviation for success reported on rollouts with each model variant.
III-C1 What is the impact of not using any auxiliary loss?
@@
We observe that the performance of our model (61%) is significantly improved by enforcing representation constraints through auxiliary losses, in comparison to using only action prediction loss (45%). It highlights the importance of the proposed auxiliary losses in §II-E.
III-C2 What is the impact of the auxiliary language loss?
@@
BC-Z proposed to use language representations to improve video representations for conditioning the policy. We compare our policy with another variant trained with all losses but the Video-Text CL. We observe only marginal improvement of 1-2% in success rate when using the language loss. This implies that video alignment and video contrastive loss contribute significantly towards performance improvement. Our results hope to serve as a promising evidence that effective video representations can be learned without auxiliary losses that use pre-trained language embeddings.
@@
Figure 8: Ablation for auxilliary losses used in Vid2Robot. We compare our proposed approach that has all auxiliary losses (green, left) with a variant without language contrastive loss that was originally proposed in BC-Z (orange, middle) and a version with no auxilliary losses (blue, right). More details in (§III-C)
@@
Figure 9: Failure analysis with policy rollouts. (Top) Policy predicts gripper pose and depends on the IK solver to move the arm. Sometimes, the IK solution can block the robot’s camera view. (Middle) Grasping failures happen, especially with transparent and deformable objects. (Bottom) Distractor objects as well as difference in lighting and background may cause recognition errors, where policy might perform the correct motion but with incorrect object(s).
@@ IV Related Work
Task Specifications for Robots The development of general-purpose robots hinges on effectively grounding task specifications. Videos are a dense source of information that not only provide what to do but also how to do it in physical world. Recent works have used videos for task specification [4, 23, 40]. Another line of work uses videos to learn world models to predict future visual observations [29, 26, 9, 31, 15]. While language [45, 7, 33, 34], final goal images [24, 6], and others like hand-drawn inputs [43] have been proposed as means for task specification, learning from prompt videos is complementary to these approaches and inevitable for rapid adaptation of trained polices to perform new manipulation skills at deployment.
@@
Learning from Human Demonstrations As videos of humans performing various tasks proliferate the internet, several works aim to address how to best leverage this information for robot learning. The difference in robot vs human embodiment poses a significant challenge, for which existing approaches range from translating image of a human into the robot [42] to inpainting for agent-agnostic representations [3]. Many prior works propose to leverage off-the-shelf models for hand pose estimation and contact tracking [4, 12, 37], object-centric representations [38, 20], as well as reward functions for reinforcement learning [3, 27, 42]. Other methods [32, 44, 4] cast this problem into visual representation learning to accelerate learning of downstream motor control tasks. While these modular learning solutions work well in limited datasets, these are prone to compounding error of each of its component, and thus, not efficiently scalable. End-to-end training approaches for goal-conditioned imitation learning [11, 41, 18, 13] and reinforcement learning [39, 35] are promising alternatives to these techniques, but these results have been largely limited in simulation and hindered by sim-to-real gap. In contrast, we choose to tackle this as an end-to-end large multi-task learning from human videos with real robot evaluations.
@@
Imitation via Paired Demonstrations Our setup of paired prompt videos and robot trajectory is most similar to One-Shot Visual Imitation literature. Many prior works assume access to pairs, where the first is used as the demonstration of the task to be performed, and the second as the observation of the agent. Some of the early works [16] proposed training a demonstration network via temporal convolution and neighorhood attention to condition a manipulation policy network. In more recent approaches like [11, 28, 20], paired demonstrations and observations are used to train a transformer policy, often with additional constraints like inverse dynamics prediction[11] or contrastive representation learning [28]. However, these approaches are largely evaluated in specific set of simulated tasks, and not compared on real robots. Most similar to our work is BC-Z [22] which reports evaluations with real robot tasks. While our setup is similar to some of this prior art, our model Vid2Robot couples large image encoders, cross-attention layers, and contrastive auxiliary losses to learn a manipulation policy that imitates a human showing a task.
@@ V Limitations and Future Directions
In §III, we show that our approach has improved over previous work but there is a gap in performance for video-conditioned policies. Language conditioned policies like [8] shows a higher success for known set of tasks with several hundreds of teleoperation trajectories for training. We, on the other hand, accomplish the first milestone of evaluating the video-conditioned policies in the similar setup. We discuss three limitations of our work and provide insights for future directions here.
First, we qualitatively investigate some reasons for failure of a policy rollout. In Fig 9, we illustrate and explain 3 examples showing how self occlusion, grasping errors and presence of distractors can lead to failure during any rollout. Second, we observe a significant drop in the grasping success in Fig 6. While we use robot camera observation to estimate the state and implicitly learn depth estimation, it is often incomplete when there is occlusion or when the robot gripper is out of camera view. By enhancing the state information with multimodal sensor fusion, we may improve the grasp success rate. Third, we consider carefully collected short task instruction demonstrations from three different sources as shown in §II-B, all of which are 5 to 20 seconds videos. To test our models on long horizon demonstrations or ‘in-the-wild’ videos online, we need effective pairing strategies for videos and a few corresponding robot trajectories to train the policy.
@@ VI Conclusion
We demonstrate novel methods for both data collection and modeling for video conditioned skill learning. These skills generalize to novel object configurations and more abstracted verb meanings when no immediately obvious object is visible. The skills and generality provided by our model complement other approaches to widen the set of skills that robots have access to, and to include skills not otherwise easily acquired. Future work can leverage these learned primitives to execute novel task plans. We hope our cross-object motion transfer experiments will encourage further research in transferring motion to new objects and settings for bootstrapping data collection, and enabling human-robot interaction with rapid adaptation to new skills.
TABLE III: Author Contributions
Name Data Collection Model Training Evaluations Infrastructure Leadership Paper Writing
Ayzaan Wahid
Christine W.Y. Chan
Danny Driess
Debidatta Dwibedi
Igor Gilitschenski
Maria Attarian
Nikhil J Joshi
Pannag R Sanketi
Pierre Sermanet
Quan Vuong
Stefan Welker
Vidhi Jain
Yonatan Bisk
Acknowledgments
@@
We would like to thank Yansong Pang, Grecia Salazar, Utsav Malla, Deeksha Manjunath, Jornell Quiambao, Sarah Nguyen, Sangeetha Ramesh, Tran Pham, Samuel Wan, Tomas Jackson, Jodilyn Peralta, Celeste Barajas, Elio Prado, Rochelle Dela Cruz, Alex Luong and Krista Reymann for supporting data collection via teleoperation. Special thanks to Jornell Quiambao, Grecia Salazar, Utsav Malla, Deeksha Manjunath, Sarah Nguyen, Sangeetha Ramesh, and Jaspiar Singh for evaluations on robot; Michael Ahn, Anthony Brohan and Keerthana Gopalakrishnan for policy evaluation infrastructure; Suneel Belkhale, Dorsa Sadigh, Chelsea Finn, and Sergey Levine for helpful discussions; Jonathan Tompson, and Vincent Vanhouke for thorough feedback on the writing. This work was also supported by Google Robotics Funding at Carnegie Mellon University.

### v2

@@ III-A2 How well do video-conditioned policies perform when shown a task in an unseen video?
III-B Tasks with More Rollouts
To comment on the statistical significance of our results, we conducted more trials while limiting the evaluation to two tasks, namely ‘place coke can upright’ and ‘close middle drawer’ for real robot policy evaluations, and reported mean success rate with confidence intervals. In total, we conducted 314 real robot rollouts for results reported in Table II.
TABLE II: Real Robot Evaluation of Vid2Robot and BC-Z with more trials to ascertain the statistical significance of the results.
Model place coke can upright close middle drawer Overall
BC-Z 19.4 $\pm$ 9.5% 39.2 $\pm$ 10.8% 30.2 $\pm$ 7.1%
Vid2Robot 39.1 $\pm$ 9.9% 48.7 $\pm$ 11.2% 43.4 $\pm$ 7.2%
III-C Cross-object motion transfer
Figure 7: Qualitative results for cross-object motion transfer. (Top-blue) Prompt video of placing coke can upright; (Green) Policy rollouts with a green can, chips bag, stapler and a soft toy in front of the robot. Vid2Robot infers the motion of place upright in the prompt video and applies it to other objects. The policy adheres to the prompt video by picking the green can instead of the chips bag or banana.
We trained our Vid2Robot and baseline with paired videos as discussed in Section II-B. Due to the pairing, the training data included only those scenarios where the interaction object shown in the prompt is present in the current robot observations. But what if we provided a prompt video of one object and tested on other objects? Does it make the same motion as shown in the prompt video? Interestingly, we found our model to perform learned manipulation actions on objects not seen in the prompt video. We call this emergent behavior as cross-object motion transfer.
We compare Vid2Robot with BC-Z for cross-object motion transfer ability with five prompt videos, namely, ‘knock water bottle over’, ‘pick green rice chip bag’, ‘place coke can upright’, ‘pick coke can from bottom drawer and place on counter’, and ‘place apple into top drawer’. We evaluate each case of a prompt video by placing unrelated objects in robot’s initial observation. The objects used for evaluation are ‘orange’, ‘green can’, ‘chips bag’, ‘banana’, ‘pink piggy soft toy’, ‘wrist watch’. We selected objects with different shapes, sizes, and deformability to evaluate situations requiring different grasps for success.
The evaluation setup is similar to Section III-A. Here, the evaluator sets up one of the objects for a task and records rollouts for each model. We compare two models on five tasks with six objects, so every evaluator runs $2\!\times\!5\!\times\!6\!=\!60$ rollouts. We repeat the evaluation with four raters, thus reporting results in Table III on $4\!\times\!60=240$ rollouts.
In Fig 7, we show the above experimental setup qualitatively. We use a prompt video to ‘place coke can upright’ and observe that the policy can transfer the action of ‘placing upright’ to objects, like a green can, a chips bag, a stapler, and a soft toy. The policy shows an implicit notion of learned pragmatics, by selecting green can over other objects.
[…]
TABLE III: Cross-object motion transfer success.
@@ III-C Cross-object motion transfer
III-D Ablations
We analyze our proposed approach to understand the following: (a) Can the policy learn possible interactions with the environment from state observations alone, without needing prompt videos? (b) How do the auxiliary loss functions impact the model’s performance?
III-D1 What is the impact of the prompt for task inference?
First, we motivate the importance of prompt videos with an example. Consider a Coke can and other objects on the countertop as the robot’s state observation. If a Coke can exists in the robot’s view, it is hard to infer whether the task is to “pick a Coke can”, “move a Coke can close to a chocolate bar”, “move a Coke can near an orange can,” or “knock over Coke can”. Furthermore, once the robot starts taking action, it can end up in new states, like being close to rxbar, which make it especially difficult to predict tasks from the current state only. Below, we empirically measure the success rate of a policy that does not have access to a suitable prompt video.
We evaluated the “no-prompt” case, in which both models see blank frames as input prompt videos. In this setup, we evaluated three tasks. For each task, we rolled out the policy 20 times. In total, we ran $2\times 20\times 3=120$ actual robot rollouts for this experiment. Here, the success rate is 23% for Vid2Robot and 5% for BC-Z over 20 rollouts per task per policy. We find that performance improves when we condition the policies on the prompt videos. The success rate improves from $5\%$ to $52.6\%$ for BC-Z and from $23\%$ to $54.6\%$ for Vid2Robot. (Refer Table I). This experiment underlines the importance of prompt videos for task success.
III-D2 What is the role of auxilliary losses?
@@
Second, we analyze the role of additional loss functions in the overall success rate. In Section II-E, we presented action prediction loss and three auxiliary losses. We investigate the impact of (1) not using any auxiliary loss and (2) adding auxiliary language loss. We consider the tasks similar to those described in Section III-A, 9 tasks for evaluating each policy. We have 3 model variants: the original Vid2Robot, the one without video-text contrastive loss (CL), and the one with only action prediction loss. We ask three human evaluators to run a model variant for two rollouts each. In total, we report results with $3\!\times\!3\!\times\!9\!\times\!2\!=\!162$ rollouts in Fig 8. The error bars show the standard deviation for success.
@@
What is the impact of not using any auxiliary loss? We observe that the performance of our model (61%) improves significantly by enforcing representation constraints through auxiliary losses, compared to using only action prediction loss (45%). It highlights the importance of the proposed auxiliary losses in Section II-E.
@@
What is the impact of the auxiliary language loss? BC-Z proposed to use language representations to improve video representations for conditioning the policy. We compare our policy with another variant trained with all losses but the Video-Text CL. We observe only a borderline improvement of 1-2% in the success rate when using language loss. This implies that video alignment and video contrastive loss contribute significantly towards performance improvement.
@@
Figure 8: Ablation for auxilliary losses used in Vid2Robot. We compare our proposed approach that has all auxiliary losses (green, left) with a variant without language contrastive loss that was originally proposed in BC-Z (orange, middle) and a version with no auxilliary losses (blue, right). More details in (§III-D)
@@
Figure 9: Failure analysis with policy rollouts. (Top) Policy predicts gripper pose and depends on the IK solver to move the arm. Sometimes, the IK solution can block the robot’s camera view. (Middle) Grasping failures happen, especially with transparent and deformable objects. (Bottom) Distractor objects and differences in lighting and background may cause recognition errors, where policy might perform the correct motion but with an incorrect object(s).
@@ IV-A Task Specifications for Robots
The development of general-purpose robots hinges on effectively grounding task specifications. Videos are a dense source of information that provides information about what to do and how to do it in the physical world. Recent works have used videos for task specification [4, 24, 42]. Another line of work uses videos to learn world models to predict future visual observations [31, 28, 10, 33, 16].
Recall our example of “open drawer”, “open cabinet”, and “open jar” in the §I. Video-conditioned policies like Vid2Robot are capable of doing these tasks because these policies identify the tasks of ”open jar” and ”open drawer” from visuals, unlike language-conditioned which have the same embedding of the verb ‘open’ for each task. Note that language is not an input to Vid2Robot. Therefore, the verb does not directly influence the action; this is a critical difference between Vid2Robot and existing language-conditioned robot policies. While language [49, 8, 35, 36], final goal images [25, 7], and others like hand-drawn inputs [45] have been proposed as means for task specification, learning from prompt videos is complementary to these approaches and inevitable for rapid adaptation of trained polices to perform new manipulation skills at deployment.
IV-B Learning from Human Demonstrations
@@
As videos of humans performing various tasks proliferate on the internet, several works aim to address how to leverage this information for robot learning. The difference in robot versus human embodiment poses a significant challenge, for which existing approaches range from translating the image of a human into the robot [44] to inpainting for agent-agnostic representations [3]. Many prior works propose to leverage off-the-shelf models for hand pose estimation and contact tracking [5, 13, 39], object-centric representations [40, 21], as well as reward functions for reinforcement learning [3, 29, 44].
XIRL [51], GraphIRL[26] and other RL approaches [41, 37] take a lot of time and manual effort for resetting scenes during the policy learning phase, limiting their applicability to real robots. Furthermore, RL often leads to unsafe situations with real-world robots. We compare Vid2Robot to another end-to-end behavior cloning method, BC-Z, that has been shown to scale to multiple tasks. Other methods [34, 47, 5] cast this problem into visual representation learning to accelerate learning of downstream motor control tasks. While these modular learning solutions work well in limited datasets, they are prone to compounding errors in each component and are not efficiently scalable. End-to-end training approaches for goal-conditioned imitation learning [12, 43, 19, 14] are also largely limited in simulation and hindered by sim-to-real gap. In contrast, we tackle this as an end-to-end large multi-task learning from human videos with real robot evaluations.
IV-C Imitation via Paired Demonstrations
@@
Our setup of paired prompt videos and robot trajectory is most similar to the One-Shot Visual Imitation literature. Many prior works assume access to pairs, where the first video demonstrates the task, and the second video shows the agent’s visual observations. Some of the early works [17] proposed training a demonstration network via temporal convolution and neighborhood attention to condition a manipulation policy network. In more recent approaches like [12, 30, 21], paired demonstrations and observations are used to train a transformer policy, often with additional constraints like inverse dynamics prediction[12] or contrastive representation learning [30]. MimicPlay [46] proposes hierarchical learning framework using human and robot teleoperated demos. However, evaluating these approaches is usually limited to a specific set of tasks, without any changes in background between the two videos.
BC-Z [22] is most similar to our work, which reports real robot evaluations. While our training setup has similarities with BC-Z, our model Vid2Robot couples large image encoders, cross-attention layers, and contrastive auxiliary losses to learn a manipulation policy that imitates a human showing a task. Recent approaches for self-supervised skill discovery like XSkill [48] learn skills from unpaired human and robot videos, while our approach uses text descriptions of the task to pair them explicitly. The paired human and robot videos contain different backgrounds, lighting, and object arrangements, thereby training the visual representations to be invariant to these settings, and focus on the task semantics instead.
@@ V Limitations and Future Directions
In Sec III, we show that our approach has improved over previous work but there is a gap in performance for video-conditioned policies. Below we discuss the limitations of our work and provide insights for the future.
First, we qualitatively investigate some reasons for the failure of a policy rollout. In Fig 9, we illustrate and explain three examples of how self-occlusion, grasping errors, and the presence of distractors can lead to failure during any rollout.
[…]
Third, we consider carefully collected short task instruction demonstrations from three different sources as shown in Section II-B, all of which are 5 to 20-second videos. To test our models on long-horizon demonstrations or ‘in-the-wild’ videos online, we need effective pairing strategies for videos and a few corresponding robot trajectories to train the policy.
@@ VI Conclusion
We present Vid2Robot, an end-to-end video-conditioned robot policy. Our proposed system trains on paired videos such that both videos demonstrate the same task but differ in diverse settings of lighting, background, and distractor objects. We use cross-attention (i) to learn the joint latent representations from prompt and state encodings and then (ii) to decode the action. We train the entire model for action prediction with cross-entropy loss and three auxiliary losses that encourage learning of generalizable latent representations to infer tasks directly from raw pixels to suitable actions. Vid2Robot outperforms BC-Z by over $\sim$20when prompted with human videos. Further, Vid2Robot outperforms BC-Z by $\sim$17for cross-object motion transfer; that is, if the prompt video didn’t have the exact object as the object the robot is manipulating now, the model still produces valid actions for the same verb but different objects. Cross-object motion transfer is a promising direction for further extending pragmatic transfer learning of motion to new objects. We hope Vid2Robot enables bootstrapping data collection and human-robot interaction with rapid adaptation to new skills.
Acknowledgements
@@
We would like to thank Yansong Pang, Grecia Salazar, Utsav Malla, Deeksha Manjunath, Jornell Quiambao, Sarah Nguyen, Sangeetha Ramesh, Tran Pham, Samuel Wan, Tomas Jackson, Jodilyn Peralta, Celeste Barajas, Elio Prado, Rochelle Dela Cruz, Alex Luong and Krista Reymann for supporting data collection via teleoperation. Special thanks to Jornell Quiambao, Grecia Salazar, Utsav Malla, Deeksha Manjunath, Sarah Nguyen, Sangeetha Ramesh, and Jaspiar Singh for evaluations on robot; Michael Ahn, Anthony Brohan and Keerthana Gopalakrishnan for policy evaluation infrastructure; Suneel Belkhale, Dorsa Sadigh, Chelsea Finn, and Sergey Levine for helpful discussions; Jonathan Tompson, and Vincent Vanhouke for thorough feedback on the writing. This work was performed by the first author as an intern at Google DeepMind Robotics. This work was also supported is partially funded by an unrestricted gift from Google and by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490375.
In our video provided in the supplementary materials, we show several examples of successful rollouts, and model architecture with prompt video and robot observations. We provide specific examples of task-based success, and cross-object motion transfer. Additionally, we provide qualitative results on long-horizon task composition and rollouts for tasks that are rare in our training dataset. Finally, we provide rollouts of noted failures like case of self-occlusion, grasping errors and ambiguous interpretation of the task in the prompt video.
Appendix A Training Details
We will release the model code and trained checkpoints. In Table IV, we present the detailed version of the Vid2Robot model architecture. In Table V, we present the hyperparameters used while training. We also add data augmentations to the videos while training as per the settings in Table VI. We normalize the videos as required for the pre-trained ViT model.
TABLE IV: Detailed Architecture of Vid2Robot.
Module Layer Output Size Notes
Prompt Encoder Input 16$\times$224$\times$224$\times$3 Reference Video of 16 frames
Image Encoder 16$\times$196$\times$768 Each frame passes through ViT-B/16
Reshape 3136$\times$768 Reshapes all space-time tokens to be in one dimension
Perceiver Resampler 64$\times$768 2 layers with 768 dims and 12 attention heads
State Encoder Input 8$\times$224$\times$224$\times$3 Robot Observation Video of 8 frames
Image Encoder 8$\times$196$\times$768 Each frame passes through ViT-B/16
Reshape 1568$\times$768 Reshapes all space-time tokens to be in one dimension
Perceiver Resampler 64$\times$768 2 layers with 768 dims and 12 attention heads
State Prompt Encoder Cross-attention 64$\times$768 4 layers of cross-attention with 768 dim and 8 attention heads
Transformer Layer Uses prompt tokens as keys and state tokens as queries
Action Decoder Cross-attention 11$\times$768 4 layers of cross-attention with 768 dim and 8 attention heads
Transformer Layer Uses prompt-aware state tokens as keys and action position embeddings as queries
Action Head Linear 11$\times$256 Projection of action decoder output to 256 bins
TABLE V: Hyperparameters for Vid2Robot.
Hyperparameters Values
Batch size 2048
Learning rate 8e-5
Optimizer AdamW
Num training steps 200,000
Warmup steps 2000
Image size 224
Num prompt frames 16
Num robot frames 8
Prediction Horizon 4
TABLE VI: Data augmentation for training Vid2Robot.
Hyperparameters Ranges
Height crop range (0.95, 1.)
Width crop range (0.95, 1.)
Brightness range (0.9, 1.1)
Contrast range (0.8, 1.2)
Hue range (0, 0.03)
Saturation range (0.8, 1.2)
Appendix B Dataset
We created the dataset mixture from the three sources. We collected 120k robot trajectories, 5k human trajectories, and 5k co-located human and robot trajectories. To create pairs, we sampled 3 videos as prompts per robot trajectory. This gives 360k pairs for Robot-Robot, about 15k pairs for Hindsight Human-Robot and 5k pairs for Co-located Human-Robot datasets. To ensure that all pairs are approximately sampled equally, we create the mixture proportions as 90% Robot-Robot, 5% Hindsight Human-Robot and 5% Co-located Human-Robot.
Appendix C Qualitative Results on Rare Tasks
We also test Vid2Robot on some tasks which are rare in the training set and find that video-conditioned models are also able to complete them. Some examples of these tasks are: opening glass jar, picking up green micro-fiber cloth and pulling out napkin. We show examples in Fig. 10. For the video version of these results please take a look at the supplementary video.
Figure 10: Qualitative results on rare tasks.
Appendix D Hardware setup
In all the experiments in this paper, we use mobile manipulators. These robots have a 7 degree-of-freedom arm, a two-fingered gripper, and a mobile base. As the focus of our experiments is on manipulating objects, we keep the base fixed for each episode.
We test the policy with client-server setup, where the policy is deployed as server, which on-robot client can query. The observations from the head camera of the robot are recorded on the client side, to maintain a recent history of 8 frames. Based on the average response time from the policy, we can execute predicted actions at approximately 5-7 Hz.
Appendix E Vid2Robot for Long Horizon Tasks
We show Vid2Robot can perform long horizon tasks by using multiple prompt videos showing different tasks. To do so, we chain different prompt videos together to create a long horizon task. For example, Vid2Robot can be used to prompt a robot to clean up the objects on a table into the drawer by chaining prompt videos of 3 tasks: open the drawer, place object on table into drawer, and close the drawer.
When Vid2Robot outputs terminate episode for the current prompt video, we change the prompt video to be the next sub-task without resetting the robot. In this manner, the robot can complete long horizon tasks with Vid2Robot without explicitly being trained on long horizon videos. We show successful examples in Fig. 11. For videos of the robot completing long-horizon tasks, check the summary video in the supplementary material.
Figure 11: Qualitative results showing how Vid2Robot can perform long horizon tasks with chaine prompt videos.